OpenStudy (anonymous):

multiply the polynomials. (7z+5)(z^2-2z+2)

OpenStudy (mathmate):

Write it as: 7z(z^2-2z+2)+5(z^2-2z+2) and distribute. Then collect terms in decreasing powers of z.

OpenStudy (kendricklamar2014):

\[(7z+5)(z^2-2z+2)\] \[=(7z+5)(z2+−2z+2)\] \[=(7z)(z2)+(7z)(−2z)+(7z)(2)+(5)(z2)+(5)(−2z)+(5)(2)\] \[=7z3−14z2+14z+5z2−10z+10\] \[=7z3−9z2+4z+10\]